OpenStudy (anonymous):

(3-a)(b+c), What are the factors in the multiplication problem? Help!!! Also 52; x/8=6, is 52 the solution to the prob?? My answer is no 48 is the solution.

OpenStudy (anonymous):

if i put 52 as a value of x (52)/8=6 gives 6.5 =6 and if i put 48 as x (48)/8=6 gives 6=6 it means 48 satisfy the equation x/8=6

OpenStudy (anonymous):

so 48 is the solution

OpenStudy (anonymous):

48 is the solution

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

welcome its wht i do you can give me a medal if u like the answer